OpenStudy (anonymous):

Name the algebraic property demonstrated in the example below: (2 + x) + y = 2 + (x + y) a.Commutative Property of Addition b.Associative Property of Addition c.Distributive Property d.Transitive Property

OpenStudy (anonymous):

Do you know the different algebraic properies ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

5 + 2 = 2 + 5 and 4 • 2 = 2 • 4 This demonstrates commutative property of addition and multiplication. 1 + (2 + 3) = (1 + 2) + 3 and 1 • (2 • 3) = (1 • 2) • 3 This demonstrates associative property of addition and multiplication 2(3 + 4) = 2(3) + 2(4) and 4(3 – 1) = 4(3) – 4(1) This demonstates distributive property If a = b and b = c, then a = c This demonstates transitive property

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem. Hope I helped :)